Learning Seminar: Experiments in Education

Syllabus, Spring 2003

1 General Information

Intructor’s Contact Information:

Jimmy Rising (jrising@mit.edu) Random Hall, rm. 244 phone: 5-9725

The seminar is scheduled to meet Tuesdays, 1:30 - 2:30, and Fridays, 3:00 - 4:00.

2 Seminar Goals

I want to explore experiments in education and learn more about how education and learning might be done. The Learning Seminar is going to be a reading and discussion class to address these issues.

This seminar is not to be a survey of experiments in education. In a sense, its goal is to determine how learning should happen and what kinds of contexts allow it to happen. I expect reflected disagreement on both points, and that more than one answer will be right. The point is both to find the answer that is right for each person and to suggest how future educational institutions might be designed to make those right ways of learning happen.

The class may be taken as either a 6 or 9 unit seminar. The units are broken down into two 1 hour meetings per week, and two hours of reading as preparation for each meeting. Students who choose to take the 9 unit version will be expected to put in an average of 1 hour of independent research per week and 2 hours of writing. They will turn in a 1 page proposal after 4 weeks, a 6-7 page first draft after 9 weeks, and a 10-12 page, significantly revised and improved final paper at the end of the semester.

I am hoping to involve several of the staff at ESG by inviting them to discuss their own ideas on education, as well as possibly bring in other instructors from around the institute.

3 Readings Syllabus

This seminar is interested in experiments others have done in education for what they can tell us about how learning happens, and it is concerned with how learning happens for what that can tell us about what experiments should be attempted. To this end, the class is structured by alternating between issues of learning and experiments in education. The topic on learning may not match up exactly with the corresponding experiment, but I have tried to take advantage of resonances between the two. Here are the readings for the classes:

1. How Children Fail - John Holt: excerpts on problems with schools

2. Zen and the Art of Motorcycle Maintenance - Robert Pirsig: chapter 16 (grade-less education)

3. How Children Fail - John Holt: excerpts on problems with schools

4. Maverick Colleges: Reed: A Middle Course - Ryo Takahashi

5. TBD: Deep vs. Shallow Learning Material; TBD: Scientific American article on brain learning (stimulus-response)

6. Situated Learning: Legitmate Peripheral Participation - J. Lave and E. Wenger

7. The Torch and the Firehose - Arthur Mattuck: excerpts on teaching

8. Olin College Propaganda

9. My Years at the Experimental Study Group - George Valley: excerpts on convergence, divergence, and basic ideas

10. My Years at the Experimental Study Group - George Valley: the early years

11. How Children Learn - John Holt: excerpts on teaching problem children

12. Summerhill School - Alexander Neill, et. al.

13. Types of Learning and Types of Intelligence

14. Waldorf Schools and The Home-schooling Movement (and other movements?)

15. Juggling for the Complete Klutz - John Cassidy

16. Maverick Colleges: Deep Springs: Loyalty to a Fault? - L. Jackson Newell

17. ”Dead Poets’ Society” - movie 18. Maverick Colleges: Antioch: Vision and Revision - Kerrie Naylor

19. Fields, Factories, and Workshops - Peter Kropotkin: excerpts on purpose of education 20. Maverick Colleges: Black Mountain: Meteor Among Mavericks

21. Rules for the Direction of the Mind - Rene Descartes: excerpts on learning slowly 22. Maverick Colleges: Chicago: Young Hutchins’ Dream

23. Zen and the Art of Archery: excerpts on non-individualized learning

24. Maverick Colleges: St. John’s: Back to Classics

25. Zen and the Art of Motorcycle Maintenance - Robert Pirsig: chapter 26 (gumption)

26. The Harrad Experiment - Robert Rimmer: Chapter 1, 2

The readings that start “Maverick Colleges” are from Maverick Colleges: Ten Notable Experiments in American Undergraduate Education, edited by L. Jackson Newell and Katherine C. Reynolds. We may also want to use more material from the shelves at ESG on learning (I have not done an extensive study of the materials there) and from education magazines.